# Harnessing Data and Physics for Deep Learning Phase Recovery

### v1

@@ Abstract
Phase recovery, calculating the phase of a light wave from its intensity measurements, is essential for various applications, such as coherent diffraction imaging, adaptive optics, and biomedical imaging. It enables the reconstruction of an object’s refractive index distribution or topography as well as the correction of imaging system aberrations. In recent years, deep learning has been proven to be highly effective in addressing phase recovery problems. Two main deep learning phase recovery strategies are data-driven (DD) with supervised learning mode and physics-driven (PD) with self-supervised learning mode. DD and PD achieve the same goal in different ways and lack the necessary study to reveal similarities and differences. Therefore, in this paper, we comprehensively compare these two deep learning phase recovery strategies in terms of time consumption, accuracy, generalization ability, ill-posedness adaptability, and prior capacity. What’s more, we propose a co-driven (CD) strategy of combining datasets and physics for the balance of high- and low-frequency information. The codes for DD, PD, and CD are publicly available at https://github.com/kqwang/DLPR.
@@ 1 Introduction
Since optical detectors, such as charge-coupled device sensors, can only record the intensity/amplitude but lose the phase, one has to recover the phase from the recorded intensity indirectly. And precisely because of the loss of the phase, it is ill-posed to directly calculate the phase on the object plane from the only amplitude on the measurement plane through the forward physical model. On the one hand, the phase can be iteratively retrieved from intensity measurements with prior knowledge, i.e., phase retrieval [6]. On the other hand, by incorporating additional information, this problem can be transformed into a well-posed one and solved directly, such as holography or interferometry with reference light [7, 8], Shack-Hartmann wavefront sensing with micro-lens array [9, 10], and the transport of intensity equation with multiple through-focus amplitudes [11, 12].
In recent years, deep learning, with artificial neural networks as the carrier, has brought new solutions to phase recovery, where the core idea is to train neural networks to learn the mapping relationship from intensity measurements to the light wave phase [13, 1]. As illustrated in Fig. 1, the training of neural networks can be driven either supervised by paired datasets as an implicit prior (data-driven, DD) or self-supervised by physical models as an explicit prior (physics-driven, PD) [1].
@@
The idea of PD phase recovery was first introduced by Boominathan et al. [37] in their simulation work on Fourier ptychography. Wang et al. [38] first experimentally used PD to iteratively infer the phase of a phase-only object from its diffraction image directly on an untrained/initialized neural network. Afterward, it was subsequently extended to the cases of unknown defocus distances [39], dual wavelengths [40], and complex-valued amplitude objects [41, 42]. In the quest for faster inference times, PD and a large number of intensity measurements were used for neural network pre-training [41, 42, 43, 44, 45]. Further, refinement of pre-trained neural networks by PD achieved higher accuracy with lower inference time [46, 47].
@@ 2 Principles
For physical processes that can be well modeled, such as phase recovery, PD is another available strategy. With a self-supervised learning mode, PD uses a numerical propagation $G(\cdot)$ as an explicit prior to drive the training or inference of neural networks (Fig. 3). Different from DD, which calculates the loss function in the phase domain, PD converts the network output from the phase domain to the hologram domain via numerical propagation and then calculates the loss function. This explicit prior can be utilized in three ways: untrained PD (uPD) [38], trained PD (tPD) [44], and tPD with refinement (tPDr) [46].
@@
With driving of the explicit prior $G(\cdot)$, uPD iteratively optimizes an untrained neural network $f_{\omega}(\cdot)$ to infer the phase $\hat{P}_{x}$ of an unseen object from its hologram $H_{x}$ (Fig. 3a):
@@
In tPD, the explicit prior $G(\cdot)$ is employed to train an untrained neural network $f_{\omega}(\cdot)$ with intensity-only training dataset $S_{H}=\{(H_{i}),i=1,\ldots,N\}$ as input, and then the trained neural network $f_{\omega^{\ast}}$ infers the phase $\hat{P}_{x}$ of an unseen object from its hologram $H_{x}$ (Fig. 3b):
@@
In tPDr, the trained neural network $f_{\omega^{\ast}}(\cdot)$ of tPD is iteratively fine-tuned on the hologram of the unseen object (Fig. 3c):
@@
3 Comparisons
To avoid unnecessary distraction factors, all datasets used for comparison are generated through numerical simulation based on ImageNet, LFW, and MNIST, see Appendix A. All methods use the same U-Net-based neural network, the specific structure of which is described in the Supplementary Material of Ref. 48. The implementation of the neural network is set uniformly, see Appendix B. The average peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) are used to quantify the inference accuracy.
@@ 3.1 Comparison of time consumption and accuracy
In terms of time consumption, DD, tPD, and tPDr all require pre-training before inference, thus consuming hours or even more for neural network optimization, whereas uPD performs inference for the tested sample directly on an untrained neural network. During the inference stage of DD and tPD, the hologram of the tested sample passes through the trained neural network once in one second, while the inference process for uPD and tPDr takes several minutes for iteration.
@@
Although having the same accuracy index (Table 2), the inference result of tPD shows better high-frequency detailed information while that of DD shows better low-frequency background information (Figure 4). According to the frequency principle, deep neural networks are more inclined to learn low-frequency information in the data [49]. DD learns the hologram-phase mapping relationship through the loss function in the phase domain, while PD uses numerical propagation to transfer it from the phase domain to the hologram domain. On the one hand, as shown in the white curve on the left side of Figure 4, the high-frequency phase information (steeper curve) is recorded in the diffraction fringes of the hologram which contains a more balanced high- and low-frequency information (smoother curve). This is more favorable for PD to learn those high-frequency phase information from the loss function in the hologram domain. On the other hand, the low-frequency phase causes only little contrast in the hologram, making it difficult for PD to learn low-frequency phase information, especially the plane background phase.
@@
where $\alpha$ is the weight used to control the contribution of the data-driven term and physical-driven term, which is set to 0.3. As illustrated in Figure 5, compared to the low-frequency-tendency DD and high-frequency-tendency tPD, CD takes into account both the high-frequency phase (see the blue box) and low-frequency phase (see the green box).
@@ 3.2 Comparison of generalization abilit
Overall, the dataset is the main factor affecting the generalization ability of the trained neural network. Specifically, the neural networks trained by ImageNet and LFW generally perform better on all three testing datasets, while the neural networks trained by MNIST can only infer the overall distribution of ImageNet and LFW but lack detailed information. Admittedly, MNIST itself lacks detailed information, so it is reasonable that neural networks trained with it would not be able to fully infer detailed information about ImangeNet and LFW. In this extreme case, tPD is significantly better than DD, both in terms of inference results and SSIM. As can be seen in Fig. 6, tPD infers more detailed information than DD (marked by the green arrow). Another thing worth noting is that for the case of using neural networks trained by ImageNet and LFW to infer MNIST, although the inference results of both tPD and DD appear to be ideal, the SSIM of tPD is much lower than that of DD. As can be seen from the absolute error maps (marked by the yellow arrow), the error in the background part of tPD is relatively larger than that of DD, which confirms a conclusion in Sec. 3.1 that tPD is not good at low-frequency phase information, especially the plane background phase.
@@ 3.4 Comparison of prior capacity
We illuminate the inference results and absolute error maps of four samples in Figure 9. As expected, DD infers the sample phase while removing the imaging aberration phase, while the inference result of tPD includes both the sample phase and the aberration phase. Accordingly, the SSIM of DD is much higher than that of tPD. In DD, the hologram contains unwanted aberration information, but the ground truth only contains sample information, which means that the dataset implicitly contains both the prior for phase recovery and the prior for aberration removal. As for tPD, the prior for the network training is derived from numerical propagation, which allows both the sample information and the aberration information in the hologram to be recovered. It should be noted that the results of uPD also contain the unwanted aberration phase just like that of tPD.
@@
4 Experimental tests
We compare DD, tPD, CD, and uPD(tPDr) using experimental holograms of standard phase objects. This experimental hologram with a defocus distance of 8.78mm is from an open-source dataset of Ref 50. To match the defocus distance of the experimental hologram, we use ImageNet to generate corresponding datasets for the network training. Inference results for all methods are given in Fig. 10.
@@ 4 Experimental tests
Overall, uPD and tPDr with multiple-times inferences have the best results, as seen from the neatly drawn peaks and valleys. It should be noted that due to the presence of redundant diffraction fringes at the edge of the hologram (see the green box in Fig. 10(a)), unwanted fluctuations appear in the background of the uPD and tPDr inference results (see the green arrows in Fig. 10(a)). Among the remaining one-time inference methods, the background fluctuations of the tPD results are larger (see the yellow arrows in Fig. 10(a)), while the detailed information of the DD results is weaker (see the yellow arrows in Fig. 10(b)). As a combination of DD and tPD, CD better considers detailed and background information. It should be noted that as the training dataset further expands, the neural network’s accuracy will increase accordingly.
@@ Appendix A Dataset generation
Three publicly available image datasets (ImageNet, LFW, and MNIST) are used to generate phases and amplitudes, and then the corresponding holograms at a certain propagation distance are computed via numerical propagation. The training and testing datasets contain 10,000 and 100 data, respectively. The size of all data is set to $256\times 256$. The propagation distance is set to 20 mm and 8.78 mm for the simulation comparisons and the experimental tests, respectively.
@@ Appendix B Network implementation
The Adam optimizer with an initial learning rate of 0.001 is adopted to update the weights and biases. The neural network training epoch of DD and PD is set to 100. The inference cycles of uPD and tPDr are set to 10,000 and 1000, respectively. All the neural networks are based on Pytorch (2.0.0) with Python (3.8.18). All operators run on a compute server equipped with AMD Ryzen Threadripper PRO 3955WX and NVIDIA GeForce RTX 3090.
Disclosures
The authors declare no competing interests.
Acknowledgments
The work was supported in part by the Research Grants Council of Hong Kong (GRF 17201620, GRF 17200321, RIF R7003-21).

### v2

# Deep learning phase recovery: data-driven, physics-driven, or combining both?

@@ Abstract
Phase recovery, calculating the phase of a light wave from its intensity measurements, is essential for various applications, such as coherent diffraction imaging, adaptive optics, and biomedical imaging. It enables the reconstruction of an object’s refractive index distribution or topography as well as the correction of imaging system aberrations. In recent years, deep learning has been proven to be highly effective in addressing phase recovery problems. Two most direct deep learning phase recovery strategies are data-driven (DD) with supervised learning mode and physics-driven (PD) with self-supervised learning mode. DD and PD achieve the same goal in different ways and lack the necessary study to reveal similarities and differences. Therefore, in this paper, we comprehensively compare these two deep learning phase recovery strategies in terms of time consumption, accuracy, generalization ability, ill-posedness adaptability, and prior capacity. What’s more, we propose a co-driven (CD) strategy of combining datasets and physics for the balance of high- and low-frequency information. The codes for DD, PD, and CD are publicly available at https://github.com/kqwang/DLPR.
@@ 1 Introduction
Since optical detectors, such as charge-coupled device sensors, can only record the intensity/amplitude but lose the phase, one has to recover the phase from the recorded intensity indirectly. And precisely because of the loss of the phase, it is ill-posed to directly calculate the phase on the object plane from the only amplitude on the measurement plane through the forward physical model. On the one hand, the phase can be iteratively retrieved from intensity measurements with prior knowledge, i.e., phase retrieval [6]. On the other hand, by incorporating additional information, this problem can be transformed into a well-posed one and solved directly, such as holography or interferometry with reference light [7, 8], Shack-Hartmann wavefront sensing with micro-lens array [9, 10], and the transport of intensity equation with multiple through-focus intensity images [11, 12].
In recent years, deep learning, with artificial neural networks as the carrier, has brought new solutions to phase recovery. One of the most direct ways is to train neural networks to learn the mapping relationship from intensity measurements to the light wave phase [13, 1, 14]. On one hand, the training of neural networks can be driven by paired input-label datasets as an implicit prior, called data-driven (DD) strategies (see the upper part of Fig. 1[1]. On the other hand, forward physical models can be used as an explicit prior to drive the training of neural networks with input-only datasets, called physics-driven (PD) strategies (see the lower part of Fig. 1[1]. In addition, neural networks can also indirectly participate in the process of phase recovery including pre-processing, in-processing (physics-connect-network, network-in-physics, and physics-in-network), and post-processing [1]. Compared with classic phase recovery methods that mainly rely on physical models, deep learning methods additionally introduce prior knowledge from datasets and neural network structures to improve efficiency.
@@
The idea of PD phase recovery was first introduced by Boominathan et al. [38] in their simulation work on Fourier ptychography. Wang et al. [39] first experimentally used PD to iteratively infer the phase of a phase-only object from its diffraction image directly on an untrained/initialized neural network. Afterward, it was subsequently extended to the cases of unknown defocus distances [40], dual wavelengths [41], and complex-valued amplitude objects [42, 43]. In the quest for faster inference times, PD and a large number of intensity measurements were used for neural network pre-training [42, 43, 44, 45, 46]. Further, refinement of pre-trained neural networks by PD achieved higher accuracy with lower inference time [47, 48]. It should be noted that the PD strategies mentioned here do not include methods that use random vectors or matrices as the inputs of neural networks. For the specific differences, please refer to the italicized part on page 22 of Ref. 1.
@@ 2 Principles and Methods
For physical processes that can be well modeled, such as phase recovery, PD is another available strategy. With a self-supervised learning mode, PD uses a numerical propagation $G(\cdot)$ as an explicit prior to drive the training or inference of neural networks (Fig. 3). Different from DD, which calculates the loss function in the phase domain, PD converts the network output from the phase domain to the hologram domain via numerical propagation $G(\cdot)$ and then calculates the loss function. This numerical propagation $G(\cdot)$ can be utilized to optimize the neural network in three ways: untrained PD (uPD) [39], trained PD (tPD) [45], and tPD with refinement (tPDr) [47].
@@
With driving of the numerical propagation $G(\cdot)$, uPD iteratively optimizes an initialized neural network $f_{\omega}(\cdot)$ to directly infer the phase $\hat{P}_{x}$ of an unseen object from its hologram $H_{x}$ (Fig. 3a):
@@
The most significant advantage of uPD is that it does not require any dataset to pre-process the neural network before inferences.
In tPD, the numerical propagation $G(\cdot)$ is employed to train the neural network $f_{\omega}(\cdot)$ with intensity-only training dataset $S_{H}=\{(H_{i}),i=1,\ldots,N\}$ as input, and then the trained neural network $f_{\omega^{\ast}}$ infers the phase $\hat{P}_{x}$ of an unseen object from its hologram $H_{x}$ (Fig. 3b):
@@
Comparing Eqs. 3 and 6, we can find that the working modes of tPD and DD are similar. However, due to the use of numerical propagation $G(\cdot)$, the training dataset for tPD only requires a large number of holograms without the corresponding phase as GT.
As a strategy combining uPD and tPD, tPDr iteratively fine-tunes the tPD trained neural network $f_{\omega^{\ast}}(\cdot)$ on the hologram of the unseen object (Fig. 3c):
@@
In addition, some methods use both forward physical models and data-driven neural networks for phase recovery. On the one hand, some methods first use forward physical models to recover preliminary phases from holograms and then use data-driven neural networks to either remove unwanted components [49, 50] or perform resolution enhancement [51, 52] or convert imaging modalities[53]. On the other hand, some methods use data-driven neural networks to generate holograms with different propagation distances from a hologram and then recover the phase using iterative algorithms based on forward physical models[54]. There is also an interesting way to introduce data-driven into physics-driven in the form of a generative adversarial network for phase recovery[55].
@@
3 Results and discussion
To avoid unnecessary distraction factors, all datasets used for comparison are generated through numerical simulation based on ImageNet, LFW, and MNIST, see Appendix A. ImageNet represents highly complex dense samples, LFW represents moderately complex dense samples, and MNIST represents simple sparse samples. Given its ubiquity in computational imaging, all methods use the same U-Net-based neural network, the specific structure of which is described in the Supplementary Material of Ref. 56. The implementation of the neural network is set uniformly, see Appendix B. The average peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) are used to quantify the inference accuracy.
@@ 3.1 Comparison of time consumption and accuracy
In terms of time consumption, DD, tPD, and tPDr all require pre-training before inference, thus consuming hours or even more for neural network optimization, whereas uPD performs inference for the tested sample directly on an initialized neural network. During the inference stage of DD and tPD, the hologram of the tested sample passes through the trained neural network once in one second, while the inference process for uPD and tPDr takes several minutes for iteration.
@@
Although having the same accuracy index (Table 2), the inference result of tPD shows better high-frequency detailed information while that of DD shows better low-frequency background information (Fig. 4). According to the frequency principle, deep neural networks are more inclined to learn low-frequency information in the data [57]. DD learns the hologram-phase mapping relationship through the loss function in the phase domain, while PD uses numerical propagation to transfer it from the phase domain to the hologram domain. On the one hand, as shown in the white curve on the left side of Fig. 4, the high-frequency phase information (steeper curve) is recorded in the diffraction fringes of the hologram which contains a more balanced high- and low-frequency information (smoother curve). This is more favorable for PD to learn those high-frequency phase information from the loss function in the hologram domain. On the other hand, the low-frequency phase causes only little contrast in the hologram, making it difficult for PD to learn low-frequency phase information, especially the plane background phase.
@@
where $\alpha$ is the weight used to control the contribution of the data-driven term and physical-driven term, which is set to 0.3. As illustrated in Fig. 5, compared to the low-frequency-tendency DD and high-frequency-tendency tPD, CD takes into account both the high-frequency phase (see the blue box) and low-frequency phase (see the green box). It should be noted that we only compared CD with DD and tPD since they all go through the neural network once for inference.
Interestingly, by comparing the inference results of holograms under different propagation distances (see Fig. S1 of the Supplementary Material), we find that DD has a higher tolerance for defocus distance than tPD. This is most likely due to the fact that the loss function used by tPD for the neural network training is calculated in the hologram domain, and thus it is more sensitive to changes in defocus holograms than DD. In addition, CD’s sensitivity to defocus distance is neutralized.
@@ 3.2 Comparison of generalization abilit
Overall, the dataset is the main factor affecting the generalization ability of the trained neural network. Specifically, the neural networks trained by ImageNet and LFW generally perform better on all three testing datasets, while the neural networks trained by MNIST can only infer the overall distribution of ImageNet and LFW but lack detailed information. Admittedly, MNIST itself lacks detailed information, so it is reasonable that neural networks trained with it would not be able to fully infer detailed information about ImangeNet and LFW. In this extreme case, tPD is significantly better than DD, both in terms of inference results and SSIM. As can be seen in Fig. 6, tPD infers more detailed information than DD (marked by the green arrow). Nonetheless, these results are sufficient to prove the strong generalization ability of DD and tPD, because MNIST used for training is very sparse handwritten digits with monotonous features, but the trained neural network can still do inference for the complex and feature-rich samples in ImageNet and LFW. Another thing worth noting is that for the case of using neural networks trained by ImageNet and LFW to infer MNIST, although the inference results of both tPD and DD appear to be ideal, the SSIM of tPD is much lower than that of DD. As can be seen from the absolute error maps (marked by the yellow arrow), the error in the background part of tPD is relatively larger than that of DD, which confirms a conclusion in Sec. 3.1 that tPD is not good at low-frequency phase information, especially the plane background phase.
@@ 3.4 Comparison of prior capacity
We illuminate the inference results and absolute error maps of four samples in Fig. 9. As expected, DD infers the sample phase while removing the imaging aberration phase, while the inference result of tPD includes both the sample phase and the aberration phase. Accordingly, the SSIM of DD is much higher than that of tPD. In DD, the hologram contains unwanted aberration information, but the ground truth only contains sample information, which means that the dataset implicitly contains both the prior for phase recovery and the prior for aberration removal. As for tPD, the prior for the network training is derived from numerical propagation, which allows both the sample information and the aberration information in the hologram to be recovered. It should be noted that the results of uPD also contain the unwanted aberration phase just like that of tPD.
@@
3.5 Comparison of experimental data
We compare DD, tPD, CD, and uPD(tPDr) using experimental holograms with a defocus distance of 8.78mm from an open-source dataset of Ref. 58. To match the defocus distance of the experimental hologram, we use ImageNet to generate corresponding datasets for the network training. Inference results of the standard phase object are given in Fig. 10.
@@ 3.5 Comparison of experimental data
Overall, uPD and tPDr with multiple-cycle inferences have the best results, as seen from the neatly drawn peaks and valleys. It should be noted that due to the presence of redundant diffraction fringes at the edge of the hologram (see the green box in Fig. 10(a)), unwanted fluctuations appear in the background of the uPD and tPDr inference results (see the green arrows in Fig. 10(a)). Among the remaining one-time inference methods, the background fluctuations of the tPD results are larger (see the yellow arrows in Fig. 10(a)), while the detailed information of the DD results is weaker (see the yellow arrows in Fig. 10(b)). As a combination of DD and tPD, CD better considers detailed and background information. It should be noted that as the training dataset further expands, the neural network’s accuracy will increase accordingly. In addition, we also test tissue slices and get similar conclusions, as detailed in Fig. S2 of the Supplementary Material.
@@ 4 Conclusion
We list some related papers with open-source code for readers to make further comparisons[45, 50, 47, 46].
@@ Appendix A Dataset generation
Three publicly available image datasets (ImageNet, LFW, and MNIST) are used to generate phases and amplitudes, and then the corresponding holograms at a certain propagation distance are computed via numerical propagation. The training and testing datasets contain 10,000 and 100 data, respectively. The size of all data is set to $256\times 256$. The propagation distance is set to 20 mm and 8.78 mm for the simulation comparisons and the experimental tests, respectively. In the code, we provide a hyperparameter “pad” to choose whether to use the way of “padding and cropping” to eliminate edge diffraction effects (see Fig. S3 of the Supplementary Material).
@@ Appendix B Network implementation
The Adam optimizer with an initial learning rate of 0.001 is adopted to update the weights and biases. The Adam weight decay of uPD and tPDr is set to 0.001. The learning rate decreases to 0.95 of its current value every 5 or 10 epochs until it approaches 0.00001. The batch size of DD, tPD and CD is set to 16. The neural network training epoch of DD and PD is set to 100. The inference cycles of uPD and tPDr are set to 10,000 and 1000, respectively. All the neural networks are based on Pytorch (2.0.0) with Python (3.8.18). All operators run on a compute server equipped with AMD Ryzen Threadripper PRO 3955WX and NVIDIA GeForce RTX 3090.
Code and Data, and Materials Availability
Code and data are available at https://github.com/kqwang/DLPR.
Supplementary Material for
Deep learning phase recovery: data-driven, physics-driven, or combining both?
[…]
To explore the tolerance of methods to defocus distances, we generate holograms with propagation distances from 15mm to 25mm and infer them using neural networks trained on the 20mm dataset. The SSIM of the inference results and the corresponding samples are shown in Fig. S1. It can be seen that DD is more tolerant than tPD, while CD neutralizes them.
Figure S1: Defocus distance tolerance tests of DD, tPD, and CD.
We test all methods using holograms of tissue slices. The results are shown in Fig. S2. Compared with DD, as a joint strategy of data and physics, CD also infers more high-frequency information like tPD, see the yellow arrow of Fig. S2. Since the inferences go through multiple cycles, the results of uPD and tPDr contain richer information, see the green arrow of Fig. S2.
Figure S2: Experimental tests of DD, tPD, CD, and uPD(tPDr) for tissue slices.
We provide a hyperparameter “pad” in the code to eliminate edge diffraction effects through padding and cropping. As shown in Fig. S3, a hologram is generated directly through numerical propagation (the upper part), or “padding and cropping” is added to eliminate edge diffraction effects (the lower part).
Figure S3: An example of hologram generation. Generate holograms directly by numerical propagation (upper part) or use “padding and cropping” to eliminate edge diffraction effects (lower part).